OpenStudy (anonymous):

How do you solve 6✔6+18 ______ 6

OpenStudy (anonymous):

what is that check symbol

OpenStudy (anonymous):

Square root over the 6

OpenStudy (anonymous):

\[\frac{ 6\sqrt{6}+18 }{ 6 }\]

OpenStudy (anonymous):

Try rewriting it into two separate fractions and let me know what you think then..

OpenStudy (anonymous):

I don't have square root symbol so gonna use the Lil ✔ sorry So : 6✔6 18 ___ + ___ 6 6

OpenStudy (anonymous):

What do you see now....?

OpenStudy (anonymous):

18/6 is just 3

OpenStudy (anonymous):

lets say sqrt_6 = x\[\frac{ 6x }{ 6 }\]The 6's here would just cancel and you'd be left with x. Same principle applies

OpenStudy (anonymous):

Simplified Would this be the correct answer 6✔+3

OpenStudy (anonymous):

\[\sqrt{6}+3\]Good job

OpenStudy (anonymous):

Oops that's what I meant. So another one (✔3-3)(✔5-4) Then i Foil: ✔15 -4 ✔3- 3✔5 +12 Is this it or ..?

OpenStudy (anonymous):

Looks good to me

OpenStudy (anonymous):

Is there more to it or is that the final answer

OpenStudy (anonymous):

I think that is final. I can't see how it can be simplified further.

OpenStudy (anonymous):

Omg thanks so much..got few more ✔2 ✔1 __ • __ ✔9 ✔3 My answer: ✔6 -- 3 ..??

OpenStudy (anonymous):

And: 5 __ 8-✔5 My answer : 5✔5 ___ 3

OpenStudy (anonymous):

Just wanna make sure they are correct